​(NPV with varying required rates of return​) Gubanich Sportswear is considering building a new factory to produce aluminum baseball bats. This project would require an initial cash outlay of ​$4,000,000 and would generate annual free cash inflows of ​$1,200,000 per year for 7 years. Calculate the​ project's NPV ​given:

a. A required rate of return of 9 percent

b. A required rate of return of 11 percent

c. A required rate of return of 13 percent

d. A required rate of return of 18 percent

1. If the required rate of return is 9 ​percent, the​ project's NPV is ​$
2. If the required rate of return is 11 ​percent, the​ project's NPV is ​$
3. If the required rate of return is 13 ​percent, the​ project's NPV is ​$
4. If the required rate of return is 18 ​percent, the​ project's NPV is ​$
Answer #1

a.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=1,200,000[1-(1.09)^-7]/0.09

=1,200,000*5.03295284

=6039543.40

NPV=Present value of inflows-Present value of outflows

=6039543.40-4,000,000

=$2,039,543.40(Approx).

b.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=1,200,000[1-(1.11)^-7]/0.11

=1,200,000*4.71219626

=5654635.52

NPV=Present value of inflows-Present value of outflows

=5654635.52-4,000,000

=$1,654,635.52(Approx).

c.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=1,200,000[1-(1.13)^-7]/0.13

=1,200,000*4.42261043

=5307132.52

NPV=Present value of inflows-Present value of outflows

=5307132.52-4,000,000

=$1,307,132.52(Approx).

d.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=1,200,000[1-(1.18)^-7]/0.18

=1,200,000*3.81152759

=4573833.11

NPV=Present value of inflows-Present value of outflows

=4573833.11-4,000,000

=$573,833.11(Approx).